# Similarity-Distance-Magnitude Activations

### v2

@@ Abstract
We introduce the Similarity-Distance-Magnitude (sdm) activation function, a more robust and interpretable formulation of the standard $\operatorname{softmax}$ activation function, adding Similarity (i.e., correctly predicted depth-matches into training) awareness and Distance-to-training-distribution awareness to the existing output Magnitude (i.e., decision-boundary) awareness, and enabling interpretability-by-exemplar via dense matching. We further introduce the sdm estimator, based on a data-driven partitioning of the class-wise empirical CDFs via the $\operatorname{\text{{sdm}}}$ activation, to control the class- and prediction-conditional accuracy among selective classifications. When used as the final-layer activation over pre-trained language models for selective classification, the $\operatorname{\text{{sdm}}}$ estimator is more robust to co-variate shifts and out-of-distribution inputs than existing calibration methods using $\operatorname{softmax}$ activations, while remaining informative over in-distribution data.
@@ 1 Introduction
Neural-network-based language models (LMs) pose a challenge for interpretable and reliable deployment given the non-identifiability of their parameters (Hwang & Ding, 1997, inter alia) [^1] Informally, this means that two or more distinct sets of values for the parameters can result in identical output distributions. As a consequence, interpreting the parameters of such models is typically much more complicated than with a simple linear regression model, for example., which can number in the billions or more. Instead of directly interpreting parameters, one option is to move the focus of interpretation to auditing predictions as a form of interpretability by example, or exemplar, over the representation space of such models via dense matching. However, for real-world deployments, robust approaches for predictive uncertainty are also needed, both for human decision-making and for constructing sequentially dependent LM pipelines.
@@
Known theoretical results limit the statistical quantities that can be derived over LMs. Statistical guarantees in the distribution-free setting are limited to approximately conditional quantities (Valiant, 1984; Lei & Wasserman, 2014; Foygel Barber et al., 2020, inter alia). Further, even typical approximately conditional quantities can be difficult to obtain in practice, since the minimal assumption of exchangeability with a known held-out data set is itself often violated with co-variate and label shifts, which can be difficult to foresee with existing methods. Epistemologically, the prevalence of hallucinations and highly-confident wrong answers with widely deployed LMs suggests a technical impasse in effectively modeling the predictive uncertainty, despite significant work from Bayesian, Frequentist, and empirically motivated perspectives (Gal & Ghahramani, 2016; Angelopoulos et al., 2021; Guo et al., 2017; Lakshminarayanan et al., 2017; Ovadia et al., 2019, inter alia). A foundational piece is evidently missing from the picture.
@@
Given these intrinsic challenges, we approach the problem of uncertainty quantification over LMs from a new angle and ask: Can we leverage the metric learning and dense matching capabilities of neural networks over high-dimensional inputs to at least aim to maximize, with minimal distributional assumptions, the separation of aleatoric (irreducible) uncertainty and epistemic (reducible) uncertainty, decomposing the sources of the latter in a manner that is interpretable and actionable?
@@
We answer this question in the affirmative with a conceptually parsimonious, data-driven partitioning of the data to decompose sources of epistemic uncertainty: Correctly predicted depth-matches into the training set ($\operatorname{\text{{Similarity}}}$), the $\operatorname{\text{{Distance}}}$ to the training set, and the distance to the decision-boundary ($\operatorname{\text{{Magnitude}}}$). We use these signals to construct a new activation function, the $\operatorname{\text{{sdm}}}$ activation, which can be used as a replacement for the standard $\operatorname{softmax}$ activation, as, for example, the final-layer activation. The $\operatorname{\text{{sdm}}}$ activation enables more reliable estimates of the predictive uncertainty for selective classification (Chow, 1957; Geifman & El-Yaniv, 2017, inter alia), which addresses the need for uncertainty quantification with LMs used in multi-stage decision pipelines, in settings subject to co-variate shifts and out-of-distribution inputs.
@@
  • We introduce the $\operatorname{\text{{Similarity}}}$-$\operatorname{\text{{Distance}}}$-$\operatorname{\text{{Magnitude}}}$ ($\operatorname{\text{{sdm}}}$) activation function, which encodes strong signals of epistemic uncertainty, for use with neural networks.
  • We introduce the $\operatorname{\text{{sdm}}}$ estimator for use in controlling class- and prediction-conditional accuracy among selective classifications, based on a data-driven partitioning of the class-wise empirical cumulative distribution functions (eCDFs) over the output via the $\operatorname{\text{{sdm}}}$ activation.
  • We examine the behavior of the $\operatorname{\text{{sdm}}}$ activation as a final-layer activation over pre-trained language models, using the $\operatorname{\text{{sdm}}}$ estimator for selective classification. We demonstrate empirically that the $\operatorname{\text{{sdm}}}$ estimator is more robust to co-variate-shifts and out-of-distribution inputs than existing classes of post-hoc calibration methods over $\operatorname{softmax}$ activations, while remaining informative over in-distribution data.
@@ 2.1 Setting
We consider the standard multi-class classification setting (e.g., predicting the sentiment of a movie review). We are given a training dataset, $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}=\{({\bm{x}}_{n},y_{n})\}_{n=1}^{N}$ of inputs, ${\bm{x}}\in{\mathcal{X}}$, paired with their corresponding ground-truth discrete labels, $y\in{\mathcal{Y}}=\{1,\ldots,C\}$, and a labeled calibration dataset, $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, drawn from the same distribution as $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. We are then given a new test instance, $\bm{x}$}, from an unlabeled test set, $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, and seek to estimate the label with a prediction, $\hat{y}$, via the un-normalized log probabilities (“logits”, informally) of a final linear layer: ${\bm{z}}={\bm{W}}^{T}{\bm{h}}+{\bm{b}}$, where ${\bm{h}}=\operatorname{{\rm{network}}}({\bm{x}};\theta)$ is the final hidden state of a network parameterized by $\theta$. The discrete prediction is taken as $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}}$; however, for learning $\theta$, $\bm{W}$}, and $\bm{b}$}, and for human decision-making, we also seek an estimate of the predictive uncertainty, $p(y\,|\,{\bm{x}})$, which is typically obtained by normalizing $\bm{z}$} via the $\operatorname{softmax}$ activation described next. We will make a distinction between models, $\mathcal{M}$} (defined by $\theta$, $\bm{W}$}, and $\bm{b}$}, and when applicable, the exemplar adaptor, described below), which produce the prediction, $\hat{y}$, and estimators, $\mathcal{E}$}, which provide an estimate of $p(y\,|\,{\bm{x}})$, because different estimators can be used over the same model.
@@ 3.1 Similarity-Distance-Magnitude Activation Functions
Calculating the $\operatorname{\text{{sdm}}}$ activation involves training an exemplar adaptor, a 1-D CNN adaptor (with a final linear layer) over the frozen hidden states of a network, to induce distilled, compressed representations of the underlying network’s representation space conditional on its predictions. The resulting representations provide a probabilistic mapping to the training, or support, set. In this way, neural networks, including large pre-trained networks, can be viewed as hidden instance-based metric learners, from which we can then derive signals of the epistemic uncertainty.
@@ 3.1.1 Exemplar Adaptor
We take as the CNN of our exemplar adaptor $g:{\bm{h}}\in\mathbb{R}^{D}\mapsto{\bm{h}}^{\prime}\in\mathbb{R}^{M}$, a 1-D CNN that takes as input $\bm{h}$} of the underlying network. The CNN has $M$ filters, the filter applications of which produce $\bm{h}$$^{\prime}$, the distilled representation of the underlying network. A final linear layer, ${\bm{z}}^{\prime}={\bm{W}}^{\prime T}{\bm{h}}^{\prime}+{\bm{b}}^{\prime},{\bm{z}}^{\prime}\in\mathbb{R}^{C}$, then replaces the underlying network’s linear layer, with the discrete prediction taken as $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}^{\prime}}$. This exemplar adaptor will then enable us to derive the $\operatorname{\text{{Similarity}}}$, $\operatorname{\text{{Distance}}}$, and $\operatorname{\text{{Magnitude}}}$ values, as defined next.
@@ 3.1.2 $\operatorname{\text{{Similarity}}}$
$q=\sum^{|\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}|}_{i=1}\mathbf{1}_{\hat{y}=\hat{y}^{\rm{tr}}_{(i)}}\cdot\mathbf{1}_{\hat{y}^{\rm{tr}}_{(i)}=y^{\rm{tr}}_{(i)}}\cdot\mathbf{1}_{i-1=\sum^{i-1}_{j=1}\mathbf{1}_{\hat{y}=\hat{y}^{\rm{tr}}_{(j)}}\cdot\mathbf{1}_{\hat{y}^{\rm{tr}}_{(j)}=y^{\rm{tr}}_{(j)}}}$ (4)
where the rightmost indicator function, $\mathbf{1}\in\{0,1\}$, ensures consecutive (depth-wise) matches. [^3] This seemingly simple rule differs from traditional KNN rules (Cover & Hart, 1967; Devroye et al., 1996, inter alia) in two critical respects: The neural network serves as a semi-supervised metric learner of the distances between the dense representations that identify the instances, and there is a model prediction (in addition to the ground-truth label) for each instance in the support set. The former enables effective partitioning despite the curse of high dimensions; the latter provides an additional indicator of reliability for each instance. By definition, $\operatorname{\text{$q$}}$ cannot exceed the count of the most prevalent class label in $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$, and since we assume an approximately equal number of points for each class, $\operatorname{\text{$q$}}\ll|\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}|$ is typical. For the special case of calculating $q$ for ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$, which only occurs during learning, we exclude the self-match.
@@ 3.1.3 $\operatorname{\text{{Distance}}}$
$\operatorname{\text{$d$}}=\min\left[1-{\rm{eCDF}}^{y_{1}}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}}),\ldots,1-{\rm{eCDF}}^{y_{C}}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}})\right]$ (5)
@@ 3.1.6 SDM Activation: Loss and Training
${\mathcal{L}}({\bm{G}},{\bm{W}}^{\prime},{\bm{b}}^{\prime};\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}})=-\frac{1}{N}\sum_{n}^{N}\log_{(2+q)}\left(\frac{(2+q)^{\operatorname{\text{$d$}}\cdot z^{\prime}_{y_{n}}}}{\sum^{C}_{c=1}{(2+q)^{\operatorname{\text{$d$}}\cdot z^{\prime}_{c}}}}\right)$ (7)
@@ 3.2 Evaluating Selective Classification
Quantity I corresponds to top-label calibration (Gupta & Ramdas, 2022), but with a single bin for evaluation, $\alpha,1]$, removing ambiguity with regard to the choice of binning the probabilities. Quantity II does not directly correspond to quantities typically examined in the calibration literature (Brier, 1950; Dawid, 1982; Guo et al., 2017; Vaicenavicius et al., 2019; Kull et al., 2019, inter alia), but it approximates [^4] For simplicity of presentation, we omit consideration of the Beta-distributed error term that is a function of the effective sample size of split-conformal coverage (Vovk, 2012). In practice, this term is negligible in the present experiments given $|\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}|$ and the resolution of the comparisons. See Appendix A.3 for a further discussion of analyzing the effective sample size for both the class-conditional and prediction-conditional estimates. label-conditional conformal coverage in the special case of class-wise thresholds that only admit prediction sets of cardinality 1. We introduce a straightforward procedure to estimate this quantity next.
@@ 3.2.1 Controlling the Class-conditional Accuracy among Selective Classifications with SDM Estimators
In general, the statistical coverage guarantee of marginal split-conformal estimators is not directly informative for selection, since the coverage guarantee is not conditional on the set size. We may instead seek one of various approximately conditional notions of coverage (Romano et al., 2020; Angelopoulos et al., 2021, inter alia); however, there is no guarantee that when we stratify by sets of cardinality 1, coverage will be maintained. However, there is a special case in which label-conditional conformal estimators do provide a meaningful notion of class-conditional coverage for selection. Assuming $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ are exchangeable, if the conformity score for each label is from a categorical distribution and the resulting thresholding of the class-wise empirical CDFs results in class-wise thresholds that are all greater than $\frac{1}{C}$, then the cardinality 1 sets will, on average, obtain class-conditional coverage, by definition. [^5] Although the formal interpretations are not identical, the evaluation of class-conditional coverage for a single estimate of such cardinality 1 prediction sets in this restricted setting over $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ is numerically equivalent to assessing the class-conditional accuracy, when not considering the sample size error term. Unfortunately, it may be rare to encounter this restricted setting over the full data distributions of real-world tasks. Instead, we will use the $\operatorname{\text{{sdm}}}$ activation to estimate a partitioning of the distribution into a region that approximately fulfills these assumptions.
First, we rescale the $\operatorname{\text{{Similarity}}}$ estimate to take into account the $\operatorname{\text{{Distance}}}$ and $\operatorname{\text{{Magnitude}}}$, given the predicted class. The resulting value [^6] The $\min$ ensures that $q=0$ instances remain 0. will be the basis for partitioning the distribution:
@@
where $\bot$ indicates a rejected (non-admitted) point and $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}^{\prime}}$.
To calculate this quantity for new, unseen test points ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, we require $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ to calculate $\operatorname{\text{$q$}}$ and $\operatorname{\text{$d_{\rm{nearest}}$}}$; the cached class-wise empirical CDFs of the distances over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ of Eq. 5; and $q^{\prime}_{\rm{min}}$ and the thresholds, $\psi_{1},\ldots\psi_{C}]$. Evaluation of the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ selection criteria is straightforward: We simply assess the conditional accuracies for the admitted points after stratifying by the predictions and the true labels, each in turn.
@@
The convention is to refer to the basic architecture of Eq. 6 as the $\operatorname{\text{{sdm}}}$ activation function, and using the activation with the selection criteria of Eq. 9 as the $\operatorname{\text{{sdm}}}$ estimator.

### v3

@@ Abstract
We introduce the Similarity-Distance-Magnitude (sdm) activation function, a more robust and interpretable formulation of the standard $\operatorname{softmax}$ activation function, adding Similarity (i.e., correctly predicted depth-matches into training) awareness and Distance-to-training-distribution awareness to the existing output Magnitude (i.e., decision-boundary) awareness, and enabling interpretability-by-exemplar via dense matching. We further introduce the $\operatorname{\text{{sdm}}}$ estimator, based on a data-driven partitioning of the class-wise empirical CDFs via the $\operatorname{\text{{sdm}}}$ activation, to control the class- and prediction-conditional accuracy among selective classifications. When used as the final-layer activation over pre-trained language models for selective classification, the $\operatorname{\text{{sdm}}}$ estimator is more robust to co-variate shifts and out-of-distribution inputs than existing calibration methods using $\operatorname{softmax}$ activations, while remaining informative over in-distribution data.
Similarity-Distance-Magnitude Activations
Allen Schmaltz Reexpress AI allen@re.express
@@ 1 Introduction
Neural-network-based language models (LMs) pose a challenge for interpretable and reliable deployment given the non-identifiability of their parameters (Hwang and Ding, 1997, inter alia) [^1] Informally, this means that two or more distinct sets of values for the parameters can result in identical output distributions. As a consequence, interpreting the parameters of such models is typically more complicated than with a simple linear regression model, for example., which can number in the billions or more. Instead of directly interpreting parameters, one option is to move the focus of interpretation to auditing predictions as a form of interpretability by example, or exemplar, over the representation space of such models via dense matching (Schmaltz, 2021). However, for real-world deployments, robust approaches for predictive uncertainty are also needed, both for human decision-making and for constructing sequentially dependent LM pipelines.
@@
Known theoretical results limit the statistical quantities that can be derived over LMs. Statistical guarantees in the distribution-free setting are limited to approximately conditional quantities (Valiant, 1984; Lei and Wasserman, 2014; Foygel Barber et al., 2020, inter alia). Further, even typical approximately conditional quantities can be difficult to obtain in practice, since the minimal assumption of exchangeability with a known held-out dataset is itself often violated with co-variate and label shifts, which can be difficult to foresee with existing methods. Epistemologically, the prevalence of hallucinations and highly-confident wrong answers with widely deployed LMs suggests a technical impasse in effectively modeling the predictive uncertainty, despite significant work from Bayesian, Frequentist, and empirically motivated perspectives (Gal and Ghahramani, 2016; Angelopoulos et al., 2021; Guo et al., 2017; Lakshminarayanan et al., 2017; Ovadia et al., 2019, inter alia). A foundational piece is evidently missing from the picture.
@@
Given these intrinsic challenges, we approach the problem of uncertainty quantification over LMs from a new angle and ask: Can we leverage the metric learning and dense matching capabilities of neural networks over high-dimensional inputs to at least aim to decompose signals of epistemic (reducible) uncertainty in a manner that is interpretable and actionable?
@@
We address this question with a conceptually parsimonious, data-driven partitioning of the data to decompose sources of epistemic uncertainty: Correctly predicted depth-matches into the training set ($\operatorname{\text{{Similarity}}}$), the $\operatorname{\text{{Distance}}}$ to the training set, and the distance to the decision-boundary ($\operatorname{\text{{Magnitude}}}$). We use these signals to construct a new activation function, the $\operatorname{\text{{sdm}}}$ activation, which can be used as a replacement for the standard $\operatorname{softmax}$ activation, as, for example, the final-layer activation. The $\operatorname{\text{{sdm}}}$ activation enables more reliable estimates of the predictive uncertainty for selective classification (Chow, 1957; Geifman and El-Yaniv, 2017, inter alia), which addresses the need for uncertainty quantification with LMs used in multi-stage decision pipelines, in settings subject to co-variate shifts and out-of-distribution inputs.
@@
  • We introduce the $\operatorname{\text{{Similarity}}}$-$\operatorname{\text{{Distance}}}$-$\operatorname{\text{{Magnitude}}}$ ($\operatorname{\text{{sdm}}}$) activation function.
  • We introduce the $\operatorname{\text{{sdm}}}$ estimator for use in controlling class- and prediction-conditional accuracy among selective classifications, based on a data-driven partitioning of the class-wise empirical cumulative distribution functions (eCDFs) over the $\operatorname{\text{{sdm}}}$ activation output.
  • We examine the behavior of the $\operatorname{\text{{sdm}}}$ activation as a final-layer activation over pre-trained language models, using the $\operatorname{\text{{sdm}}}$ estimator for selective classification. We demonstrate empirically that the $\operatorname{\text{{sdm}}}$ estimator is more robust to co-variate-shifts and out-of-distribution inputs than existing classes of post-hoc calibration methods, while remaining informative over in-distribution data.
@@ 2.1 Setting
We consider the standard multi-class classification setting. We are given a training dataset, $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}=\{({\bm{x}}_{n},y_{n})\}_{n=1}^{N}$ of inputs, ${\bm{x}}\in{\mathcal{X}}$, paired with their corresponding ground-truth discrete labels, $y\in{\mathcal{Y}}=\{1,\ldots,C\}$, and a labeled calibration dataset, $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, drawn from the same distribution as $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. We are then given a new test instance, $\bm{x}$}, from an unlabeled test set, $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, and seek to estimate the label with a prediction, $\hat{y}$, via the un-normalized log probabilities (“logits”, informally) of a final linear layer: ${\bm{z}}={\bm{W}}^{T}{\bm{h}}+{\bm{b}}$, where ${\bm{h}}=\operatorname{{\rm{network}}}({\bm{x}};\theta)$ is the final hidden state of a network parameterized by $\theta$. The discrete prediction is taken as $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}}$; however, for learning $\theta$, $\bm{W}$}, and $\bm{b}$}, and for human decision-making, we also seek an estimate of the predictive uncertainty, $p(y\,|\,{\bm{x}})$, which is typically obtained by normalizing $\bm{z}$} via the $\operatorname{softmax}$ activation described next. We will make a distinction between models, $\mathcal{M}$} (defined by $\theta$, $\bm{W}$}, and $\bm{b}$}, and when applicable, the exemplar adaptor, described below), which produce the prediction, $\hat{y}$, and estimators, $\mathcal{E}$}, which provide an estimate of $p(y\,|\,{\bm{x}})$, because different estimators can be used over the same model.
@@ 3.1 Similarity-Distance-Magnitude Activation Functions
Calculating the $\operatorname{\text{{sdm}}}$ activation involves training an exemplar adaptor, a 1-D CNN adaptor (with a final linear layer) over the frozen hidden states of a network, to induce distilled, compressed representations of the underlying network’s representation space conditional on its predictions. The resulting representations provide a probabilistic mapping to the training, or support, set. In this way, neural networks, including large pre-trained networks, can be viewed as hidden instance-based metric learners (Schmaltz, 2021), from which we can then derive signals of the epistemic uncertainty.
@@ 3.1.1 Exemplar Adaptor
We take as the CNN of our exemplar adaptor $g:{\bm{h}}\in\mathbb{R}^{D}\mapsto{\bm{h}}^{\prime}\in\mathbb{R}^{M}$, a 1-D CNN that takes as input $\bm{h}$} of the underlying network. The CNN has $M$ filters, the filter applications of which produce $\bm{h}$$^{\prime}$, the distilled representation of the underlying network. A final linear layer, ${\bm{z}}^{\prime}={\bm{W}}^{\prime T}{\bm{h}}^{\prime}+{\bm{b}}^{\prime},{\bm{z}}^{\prime}\in\mathbb{R}^{C}$, then replaces the underlying network’s final linear layer, with the discrete prediction taken as $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}^{\prime}}$. This exemplar adaptor will then enable us to derive the $\operatorname{\text{{Similarity}}}$, $\operatorname{\text{{Distance}}}$, and $\operatorname{\text{{Magnitude}}}$ values, as defined next.
@@ 3.1.2 $\operatorname{\text{{Similarity}}}$
$q=$
$\sum^{|\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}\rvert}_{i=1}\mathbf{1}_{\hat{y}=\hat{y}^{\rm{tr}}_{(i)}}\cdot\mathbf{1}_{\hat{y}^{\rm{tr}}_{(i)}=y^{\rm{tr}}_{(i)}}\cdot\mathbf{1}_{i-1=\sum^{i-1}_{j=1}\mathbf{1}_{\hat{y}=\hat{y}^{\rm{tr}}_{(j)}}\cdot\mathbf{1}_{\hat{y}^{\rm{tr}}_{(j)}=y^{\rm{tr}}_{(j)}}}$ (4)
where the rightmost indicator function, $\mathbf{1}\in\{0,1\}$, ensures consecutive (depth-wise) matches. [^3] This seemingly simple rule differs from traditional KNN rules (Cover and Hart, 1967; Devroye et al., 1996, inter alia) in two critical respects: The neural network serves as a semi-supervised metric learner of the distances between the dense representations that identify the instances, and there is a model prediction (in addition to the ground-truth label) for each instance in the support set. The former enables effective partitioning despite the curse of high dimensions; the latter provides an additional indicator of reliability for each instance. By definition, $\operatorname{\text{$q$}}$ cannot exceed the count of the most prevalent class label in $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$, and since we assume an approximately equal number of points for each class, $\operatorname{\text{$q$}}\ll|\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}|$ is typical. For the special case of calculating $q$ for ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$, which only occurs during learning, we exclude the self-match.
@@ 3.1.3 $\operatorname{\text{{Distance}}}$
$\operatorname{\text{$d$}}=\min$~ $[1-{\rm{eCDF}}^{y_{1}}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}}),\ldots,$
$~1-{\rm{eCDF}}^{y_{C}}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}})]$ (5)
@@ 3.1.6 SDM Activation: Loss and Training
${\mathcal{L}}({\bm{G}},{\bm{W}}^{\prime},{\bm{b}}^{\prime};\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}})=$
$-\frac{1}{N}\sum_{n}^{N}\log_{(2+q)}\left(\frac{(2+q)^{\operatorname{\text{$d$}}\cdot z^{\prime}_{y_{n}}}}{\sum^{C}_{c=1}{(2+q)^{\operatorname{\text{$d$}}\cdot z^{\prime}_{c}}}}\right)$ (7)
@@ 3.2 Evaluating Selective Classification
Quantity I corresponds to top-label calibration (Gupta and Ramdas, 2022), but with a single bin for evaluation, $\alpha,1]$, removing ambiguity with regard to the choice of binning the probabilities. Quantity II does not directly correspond to quantities typically examined in the calibration literature (Brier, 1950; Dawid, 1982; Guo et al., 2017; Vaicenavicius et al., 2019; Kull et al., 2019, inter alia), but it approximates label-conditional conformal coverage in the special case of class-wise thresholds that only admit prediction sets of cardinality 1. We introduce a straightforward procedure to estimate this quantity next.
@@ 3.2.1 Controlling the Class-conditional Accuracy among Selective Classifications with SDM Estimators
In general, the statistical coverage guarantee of marginal split-conformal estimators is not directly informative for selection, since the coverage guarantee is not conditional on the set size. We may instead seek one of various approximately conditional notions of coverage (Romano et al., 2020; Angelopoulos et al., 2021, inter alia); however, there is no guarantee that when we stratify by sets of cardinality 1, coverage will be maintained. However, there is a special case in which label-conditional conformal estimators do provide a meaningful notion of class-conditional coverage for selection. Assuming $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ are exchangeable, if the conformity score for each label is from a categorical distribution and the resulting thresholding of the class-wise empirical CDFs results in class-wise thresholds that are all greater than $\frac{1}{C}$, then the cardinality 1 sets will, on average, obtain class-conditional coverage, by definition. [^4] Although the formal interpretations are not identical, the evaluation of class-conditional coverage for a single estimate of such cardinality 1 prediction sets in this restricted setting over $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ is numerically equivalent to assessing the class-conditional accuracy, when not considering the sample-size dependent error term. See Appendix A.6 for a further discussion of analyzing the effective sample size. Unfortunately, it may be rare to encounter this restricted setting over the full data distributions of real-world tasks. Instead, we will use the $\operatorname{\text{{sdm}}}$ activation to estimate a partitioning of the distribution into a region that approximately fulfills these assumptions.
First, we rescale the $\operatorname{\text{{Similarity}}}$ estimate to take into account the $\operatorname{\text{{Distance}}}$ and $\operatorname{\text{{Magnitude}}}$, given the predicted class. The resulting value [^5] The $\min$ ensures that $q^{\prime}$ remains 0 for points with $q=0$, which are effectively out-of-distribution points. will be the basis for partitioning the distribution:
@@
where $\bot$ indicates a rejected (non-admitted) point and $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}^{\prime}}$. [^6] The convention is to refer to the basic architecture of Eq. 6 as the $\operatorname{\text{{sdm}}}$ activation function, and using the activation with the selection criteria of Eq. 9 as the $\operatorname{\text{{sdm}}}$ estimator.
To calculate this quantity for new, unseen test points ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, we require $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ to calculate $\operatorname{\text{$q$}}$ and $\operatorname{\text{$d_{\rm{nearest}}$}}$; the cached class-wise empirical CDFs of the distances over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ of Eq. 3.1.3; and $q^{\prime}_{\rm{min}}$ and the thresholds, $\psi_{1},\ldots\psi_{C}]$. Evaluation of the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ selection criteria is straightforward: We simply assess the conditional accuracies for the admitted points after stratifying by the predictions and the true labels, each in turn.